OpenStudy (anonymous):

What is the slope of the line passing through the point (3, −4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction.

OpenStudy (zzr0ck3r):

perpendicular says take the negative multiplicative inverse of the slope and that is your new slope so we know the slope is (-1/3) so we have y = (-1/3)x+b and we are given (3,-4) so plug them in -4=(-1/3)(3)+b -4 = -1+b b = -3 so whats the equation?

OpenStudy (zzr0ck3r):

O it did not even need the eqation, just the slope. so -1/3